OpenStudy (anonymous):

Show that the equation represents a circle by rewriting it in standard form, and find the center and radius of the circle. 5x^2 + 5y^2 + 10x − y = 0

OpenStudy (anonymous):

Standard form for a circle is x^2 + y^2 == r^2.

myininaya (myininaya):

\[5x^2+10x+5y^2-y=0\] i just reordered the terms \[5(x^2+2x)+5(y^2-\frac{1}{5}y)=0\] \[(x^2+2x)+(y^2-\frac{1}{5}y)=0\] i divided by sides by 5 \[(x^2+2x+(\frac{2}{2})^2)+(y^2-\frac{1}{5}y+(\frac{1}{5 \cdot 2})^2)=(\frac{2}{2})^2+(\frac{1}{5 \cdot 2})^2\] now we have \[(x^2+2x+1^2)+(y^2-\frac{1}{5}y+(\frac{1}{10})^2)=1+\frac{1}{100}\] \[(x+1)^2+(y-\frac{1}{10})^2=\frac{101}{100}\]

myininaya (myininaya):

center is \[(-1,\frac{1}{10})\] radius is \[\sqrt{\frac{101}{100}}=\frac{\sqrt{101}}{10}\]

OpenStudy (anonymous):

x^2+y^2+2x-y/5=0 => (x+1)^2 -1 +(y-1/10)^2 -1/100=0 =>(x+1)^2 +(y-1/10)^2=101/100 this standard form of eqn of this circle..where center is (−1,1/100) and radius =sqrt(101)/10